# Manify: A Python Library for Learning Non-Euclidean Representations

### v1

@@ Abstract
We present Manify, an open-source Python library for non-Euclidean representation learning. Leveraging manifold learning techniques, Manify provides tools for learning embeddings in (products of) non-Euclidean spaces, performing classification and regression with data that lives in such spaces, and estimating the curvature of a manifold. Manify aims to advance research and applications in machine learning by offering a comprehensive suite of tools for manifold-based data analysis. Our source code, examples, datasets, results, and documentation are available at https://github.com/pchlenski/manify.
@@ 3 Library structure
manifymanifoldspredictorscurvature_estimationembeddersutils $\delta$-hyperbolicity, sectional curvature, etc. Product space GCNs, decision trees, perceptrons, SVMs Embedding via coordinate learning, VAEs, or Siamese nets Loading data, benchmarking, visuals, and other conveniences
@@ 3.1 Manifolds
3.2 Embedders
Manify.embedders is designed to generate graph and data embeddings in multiple ways. Its implementation details can be found in Appendix C. Here are the core features in embedders:
@@ 3.2 Embedders
  • Manify.embedder.siamese: Siamese Neural Networks compatible with product space geometry, modeled off of Corso et al. (2021).
@@
3.3 Predictors
Manify.predictors implements predictors that work on data with product manifold-valued features. All predictors follow Scikit-Learn API conventions such as including fit and predict methods. Further mathematical details on these algorithms can be found in Appendix E. Here are the core features of predictors:
@@ 3.3 Predictors
3.4 Curvature Estimation
Manify.curvature_estimation provides functions for estimating the curvature of a dataset based on a graph or matrix of pairwise distances, helping users estimate what manifolds are appropriate for embedding their data. More details are included in Appendix F.
@@ 3.4 Curvature Estimation
3.5 Utilities
Manify.utils also provides a set of convenience functions supporting typical data analysis workflows:
@@ 4 Example Usage
import torch
from manify.manifolds import ProductManifold
from manify.embedders import coordinate_learning
from manify.predictors.decision_tree import ProductSpaceDT
from manify.utils import dataloaders
@@
# Load graph data
dists, labels, _ = dataloaders.load("polblogs")
# Create product manifold
pm = ProductManifold(signature=[(1, 4)]) # S^4_1
@@
X, _ = coordinate_learning.train_coords(
pm=pm,
dists=dists / dists.max(),
burn_in_iterations=1000,
training_iterations=9000
@@
X_train, X_test, y_train, y_test = train_test_split(X, graph_labels)
tree = ProductSpaceDT(pm=pm, max_depth=3)
tree.fit(X_train, y_train)
print(tree.score(X_test, y_test))
@@ 5 Conclusion and Future Work
We present Manify, an open-source Python package for machine learning in mixed-curvature product manifolds. Although it is particularly well-suited for features (or distances) conforming to spaces of heterogeneous curvature, such as hierarchical data, networks, and certain types of biological data, it nonetheless provides the tools for carrying out machine learning methods in any of the geometries described.
Future work should include:
  • Carefully validating additional models against their references (VAEs, Siamese Neural Nets, Product Space Perceptrons, Product Space SVMs, and sectional curvature estimation);
  • Adding additional models (Transformers/Attention (Cho et al., 2023) and PCA (Tabaghi et al., 2024));
  • Fully aligning Manify with Scikit-Learn conventions; and, of course
  • Applying Manify to more datasets and workflows!
@@
All acknowledgements go at the end of the paper before appendices and references. Moreover, you are required to declare funding (financial activities supporting the submitted work) and competing interests (related financial activities outside the submitted work). More information about this disclosure can be found on the JMLR website.
@@ Appendix A Code Referenced in this Work
Table 1: This table enumerates and cites any online code resources we consulted during the creation of Manify
@@ D.1 Coordinate Learning
$\mathcal{L}_{\text{distortion}}(\mathbf{X},\mathbf{D})=\sum_{1\leq i\leq j\leq n}\left|\left(\frac{\delta_{\mathcal{P}}(x_{i},x_{i})}{D_{i,j}}\right)^{2}-1\right|$ (47)
@@ Appendix E Mathematical Details: Predictors
Manify offers a menagerie of predictors, supporting classification and regression (and occasionally link prediction) on product manifold-valued datasets. We follow Scikit-Learn (Pedregosa et al., 2011) conventions in our API design for our predictors. In particular, we ensure that each predictor has .fit(), .predict(), and (for classifiers) .predict_proba() methods available.
@@
  • Graph Convolutional Networks, as in the $\kappa$-GCN described in Bachmann et al. (2020). By ablating certain components of the $\kappa$-GCN, we can additionally create Mulilayer Perceptrons and Logistic/Linear Regression.
@@ E.2 Product Space Graph Convolutional Networks
We implemented the $\kappa$-GCN algorithm in Manify.predictors.kappa_gcn from
Bachmann et al. (2020), who introduces gyrovector operation that extends existing frameworks (Chami et al., 2019) that are limited to negative curvature to both negative and positive values that interpolates smoothly. We start by introducing these concepts.
@@
Although is not explicit in Bachmann et al. (2020) how the logits aggregate across different product manifolds, we follow Cho et al. (2023) and later Chlenski et al. (2025) in aggregating logits as the $\ell_{2}$-norm of component manifold logits, multiplied by the sign of the sum of the component inner products:
@@
Figure 3: We diagram the relationship between 9 different models, demonstrating how they are all special cases of $\kappa-GCN$. In particular, the horizontal axis tracks the relationship $\mathbb{E}^{\subset}\mathcal{M}\subset\mathcal{P}$, whereas the vertical axis tracks the relationship $\operatorname{MLR}\subset\operatorname{MLP}\subset\operatorname{GCN}$. We use “MLR” to refer generically to all one-layer models.
@@ Appendix F Mathematical Details: Curvature Estimation
Given a set of distance, one might wish to ask what the curvature of the underlying space is. We implement two popular approaches to answering this question:
@@ F.2 $\delta$-hyperbolicity
Appendix G Experiments
@@
G.1 Validating $\delta$-Hyperbolicity Estimation (Khrulkov et al. (2020))
@@
G.1.1 Validating Coordinate Learning (Gu et al 2018)

### v2

@@ Abstract
We present Manify, an open-source Python library for non-Euclidean representation learning. Leveraging manifold learning techniques, Manify provides tools for learning embeddings in (products of) non-Euclidean spaces, performing classification and regression with data that lives in such spaces, estimating the curvature of a manifold, and more. Manify aims to advance research and applications in machine learning by offering a comprehensive suite of tools for manifold-based data analysis. Our source code, examples, and documentation are available at https://github.com/pchlenski/manify.
@@ 3 Library Structure
manifymanifoldscurvature_estimationpredictorsembeddersclusteringoptimizersutils $\delta$-hyperbolicity, sectional curvature, etc. Product space GCNs, decision trees, perceptrons, SVMs Embedding via coordinate learning, VAEs, or Siamese nets Clustering algorithms on product manifolds Riemannian optimization algorithms Loading data, benchmarking, visuals, and other conveniences
@@
Manify.embedders implements methods to embed points into product manifolds on the basis of pairwise distances and/or features. All embedders follow Scikit-Learn transformer API conventions, exposing fit and transform methods. Its implementation details can be found in Appendix D. Here are the core features in embedders:
@@
  • Manify.embedders.siamese: Siamese Neural Networks compatible with product space geometry, modeled off of Corso et al. (2021).
@@
Manify.predictors implements predictors that work on data with product manifold-valued features. All predictors follow Scikit-Learn predictor API conventions, exposing fit, predict_proba, predict, and score methods. Further mathematical details on these algorithms can be found in Appendix E. Here are the core features of predictors:
@@
  • Manify.predictors.nn: Submodule containing neural network layers, including $\kappa$-GCN layers, stereographic logits, Fermi-Dirac decoder layers, and stereographic transformer blocks (Cho et al., 2023).
Manify.curvature_estimation provides functions for estimating the curvature of a dataset based on a graph or matrix of pairwise distances, helping users estimate what manifolds are appropriate for embedding their data. More details are included in Appendix F. Here are the core features of curvature_estimation:
@@
  • Manify.curvature_estimation.greedy_method: the Greedy heuristic used to construct near-optimal signatures in Tabaghi et al. (2021)
Manify.clustering provides functions for clustering points on product manifolds. Currently, it only implements the Riemannian fuzzy $K$-means algorithm proposed by Yuan et al. (2025). Manify.optimizers contains Riemannian optimizers that are not presently implemented in Geoopt. Currently, it contains an implementation of Radan, a Riemannian variant of Adan (Xie et al., 2024). Radan is also originally proposed in Yuan et al. (2025). Details for both clustering and optimizers, and their relationship to the Riemannian Fuzzy $K$-Means algorithm, are given in Appendix G.
Manify.utils also provides a set of convenience functions supporting typical use cases:
@@ 4 Example Usage
import manify
from manify.utils.dataloaders import load_hf
@@
# Load Polblogs graph from HuggingFace
features, dists, adj, labels = load_hf("polblogs")
# Create an S^4 x H^4 product manifold
pm = manify.ProductManifold(signature=[(1.0, 4), (-1.0, 4)])
@@
embedder = manify.CoordinateLearning(pm=pm)
X_embedded = embedder.fit_transform(
X=None, D=dists, burn_in_iterations=200, training_iterations=800
@@
X_train, X_test, y_train, y_test = train_test_split(X_embedded, labels)
model = manify.ProductSpaceDT(pm=pm, max_depth=3, task="classification")
model.fit(X_train, y_train)
print(model.score(X_test, y_test))
@@ 5 Conclusion and Future Work
We present Manify, an open-source Python package for machine learning in mixed-curvature product manifolds. Although it is particularly well-suited for features (or distances) conforming to spaces of heterogeneous curvature, such as hierarchical data, networks, and certain types of biological data, it nonetheless provides the tools for carrying out machine learning methods in any of the geometries described. Future work should include:
  • Carefully validating additional models against the results reported in the original papers (VAEs, Siamese Neural Nets, Product Space Perceptrons, Product Space SVMs, and sectional curvature estimation);
  • Adding additional models such as product space PCA (Tabaghi et al., 2024));
  • Fully aligning Manify with Scikit-Learn conventions;
  • Applying Manify to more datasets and workflows.
@@
This work was funded by NSF GRFP grant DGE-2036197. We thank Jinghui Yuan for contributing the Riemannian Fuzzy K-Means algorithm to the Manify library.
@@ Appendix A Code Referenced in this Work
Table 1: This table enumerates and cites any online code resources we consulted during the creation of Manify.
@@
https://github.com/sail-sg/Adan Adan Xie et al. (2024)
https://github.com/Yuan-Jinghui/Riemannian-Fuzzy-K-Means Riemannian Fuzzy $K$-Means Yuan et al. (2025)
@@ D.1 Coordinate Learning
$\mathcal{L}_{\text{distortion}}(\mathbf{X},\mathbf{D})=\sum_{1\leq i\leq j\leq n}\left|\left(\frac{\delta_{\mathcal{P}}(\mathbf{x}_{i},\mathbf{x}_{j})}{D_{i,j}}\right)^{2}-1\right|$ (47)
@@ Appendix E Mathematical Details: Predictors
Manify offers a menagerie of predictors, supporting classification and regression (and occasionally link prediction) on product manifold-valued datasets. We follow Scikit-Learn (Pedregosa et al., 2011) conventions in our API design for our predictors. In particular, we ensure that each predictor has .fit(), .predict(), .score(), and (for classifiers) .predict_proba() methods available.
@@
  • Graph Convolutional Networks, as in the $\kappa$-GCN described in Bachmann et al. (2020). By ablating certain components of the $\kappa$-GCN, we can additionally create Multi-Layer Perceptrons and Logistic/Linear Regression.
@@ E.2 Product Space Graph Convolutional Networks
We implemented the $\kappa$-GCN algorithm in Manify.predictors.kappa_gcn from Bachmann et al. (2020), who introduces gyrovector operation that extends existing frameworks (Chami et al., 2019) that are limited to negative curvature to both negative and positive values that interpolates smoothly. We start by introducing these concepts.
@@
Although it is not explicit in Bachmann et al. (2020) how the logits aggregate across different product manifolds, we follow Cho et al. (2023) and later Chlenski et al. (2025) in aggregating logits as the $\ell_{2}$-norm of component manifold logits, multiplied by the sign of the sum of the component inner products:
@@
Figure 3: We diagram the relationship between 9 different models, demonstrating how they are all special cases of $\kappa$-GCN. In particular, the horizontal axis tracks the relationship $\mathbb{E}\subset\mathcal{M}{}\subset\mathcal{P}{}$, whereas the vertical axis tracks the relationship $\operatorname{MLR}\subset\operatorname{MLP}\subset\operatorname{GCN}$. We use “MLR” to refer generically to all one-layer models.
@@ E.4 Product Space Support Vector Machines
E.5 Mixed-Curvature Stereographic Transformers
We implemented the Fully Product-Stereographic Transformer (FPS-T) algorithm in Manify.predictors.nn.layers following the approach proposed by Cho et al. (2023). The implementation is still preliminary and has not been rigorously validated. FPS-T extends the attention mechanism of standard transformers to operate on mixed-curvature spaces using the stereographic model. Each attention head operates on a stereographic space with learnable curvature that can represent Euclidean, hyperbolic, and spherical geometries.
Given a sequence of $n$ product-stereographic embeddings, the stereographic multi-head attention mechanism first computes queries ($\mathbf{Q}$), keys ($\mathbf{K}$), and values ($\mathbf{V}$) by mapping each stereographic embedding to the tangent space of the values. This key insight enables the geometric operations necessary to compute attention scores in stereographic space. Specifically, the attention scores between $\mathit{i}$-th query $\mathbf{Q}_{i}$ and $\mathit{j}$-th key $\mathbf{K}_{i}$ are computed by parallel transporting the query and key vectors to the origin and taking their Riemannian inner product:
$\alpha_{ij}$ $=\langle\text{PT}_{V_{i}\to 0}(\mathbf{Q}_{i}),\text{PT}_{V_{j}\to 0}(\mathbf{K}_{j})\rangle$ (93)
where $\text{PT}_{V_{i}\to 0}$ denotes parallel transport from point $V_{i}$ to the origin, which preserves the geometric structure during the transport operation. The values are then aggregated using the Einstein midpoint, which is the geometric generalization of weighted averaging in curved spaces:
$\text{Agg}_{\kappa}(\mathbf{V},\boldsymbol{\alpha})_{i}$ $=\frac{1}{2}\oplus_{\kappa}\left[\sum_{j=1}^{n}\alpha_{ij}\frac{\lambda_{\kappa}^{\mathbf{V_{j}}}}{\sum_{k=1}^{n}\alpha_{ik}(\lambda_{\kappa}^{\mathbf{V_{k}}}-1)}\mathbf{V}_{j}\right]$ (94)
where $\oplus_{\kappa}$ is the Möbius addition operation, and $\lambda_{\kappa}^{x}=\frac{4}{1+\kappa\|x\|^{2}}$ is the conformal factor that accounts for the curvature at point $x$. The final multi-head attention output concatenates results from all $H$ attention heads:
$\text{MHA}_{\otimes^{\kappa}}(\mathbf{X})=\big{\|}_{h=1}^{H}\text{Aggregate}_{\kappa_{h}}(\mathbf{V}^{h},\mathbf{\alpha}^{h})$ (95)
The model learns curvatures end-to-end through gradient descent, automatically adapting the geometry to match the input graph structure.
@@ Appendix F Mathematical Details: Curvature Estimation
Given a set of distances, one might wish to ask what the curvature of the underlying space is. We implement two popular approaches to answering this question:
@@
  • Greedy heuristics, in Manify.greedy_method, iteratively select submanifolds which minimize some objective function at each step.
@@ F.2 $\delta$-hyperbolicity
F.3 Greedy signature selection
We follow Tabaghi et al. (2021) in our formulation of greedy signature selection. The core idea of this algorithm is that, given some (black-box) pipeline which returns a scalar, we can iteratively build up a near-optimal product manifold $\mathcal{P}$ from a set of candidate submanifolds (by default, $\{\mathbb{H}^{2}_{-1},\ \mathbb{E}^{2},\ \mathbb{S}^{2}_{1}\}$) for a set number of iterations. At each iteration, whichever submanifold achieves the lowest loss in the pipeline gets added to the optimal signature, and the process repeats; if none of the submanifolds beat the baseline manifold, the algorithm terminates.
Although users are encouraged to write their own pipelines for greedy signature selection, we provide two pipelines as a starting point in Manify.curvature_estimation._pipelines:
  • The distortion-minimization pipeline is implemented in Manify.curvature_estimation._pipelines.distortion_pipeline, and allows users to find a signature that minimizes distortion for a given matrix of distances $\mathbf{D}$ when embedded using the Coordinate Learning embedder.
  • The predictor score-maximization pipeline is implemented in Manify.curvature_estimation._pipelines.predictor_pipeline, and handles embedding and classifying (or regressing on) graph-based data. Given a matrix of pairwise distances $\mathbf{D}$ and labels $\mathbf{y}$, this pipeline generates embeddings, trains a classifier on the embeddings, and optimizes the accuracy (or MSE, for regression) of the predictor.
Appendix G Mathematical Details: Clustering and Optimizers
This appendix provides a brief mathematical summary of the algorithms contributed by Yuan et al. (2025) and integrated into the Manifylibrary:
  • Riemannian Fuzzy $K$-Means is implemented in Manify.clustering.fuzzy_kmeans, and provides an efficient variant of $K$-means clustering designed to work with product manifolds, and
  • Riemannian Adan is implemented in Manify.optimizers.radan, and provides a Riemannian variant of the Adan algorithm. It is used to speed up the Riemannian Fuzzy $K$-Means algorithm.
G.1 Riemannian Fuzzy $K$-Means (RFK)
The Riemannian Fuzzy $K$-Means (RFK) algorithm is an efficient clustering method for data residing on a Riemannian manifold $\mathcal{M}$. It overcomes the significant computational hurdles associated with a naive extension of the standard Fuzzy $K$-Means (also known as Fuzzy $c$-Means) algorithm.
G.1.1 The Naive Extension and its Limitations
Unlike standard $K$-means clustering, Fuzzy $K$-means allows for “fuzzy,” i.e. partial, cluster assignments for each point. The objective function for standard Fuzzy $K$-Means in Euclidean space is:
$J_{m}=\sum_{i=1}^{N}\sum_{j=1}^{C}u_{ij}^{m}\|x_{i}-c_{j}\|^{2},$ (103)
where $\{x_{i}\}_{i=1}^{N}$ are the data points, $\{c_{j}\}_{j=1}^{C}$ are the cluster centers, $u_{ij}\in[0,1]$ is the fuzzy membership of point $x_{i}$ in cluster $c_{j}$, and the exponent $m>1$ controls the “fuzziness” of class assignments.
The naive extension to a Riemannian manifold (Naive Fuzzy $K$-Means, or NFK) replaces the Euclidean distance with the geodesic distance $\delta_{\mathcal{M}}(\cdot,\cdot)$ on the manifold $\mathcal{M}$:
$J_{m}(\mathbf{U},\mathbf{C})=\sum_{i=1}^{N}\sum_{j=1}^{C}u_{ij}^{m}\delta_{\mathcal{M}}(\mathbf{x_{i}},\mathbf{c_{j}})^{2},\quad\text{s.t.}\quad\mathbf{c}_{j}\in\mathcal{M},\ \sum_{j=1}^{C}u_{ij}=1.$ (104)
While the update for the membership matrix $\mathbf{U}=[u_{ij}]$ retains a closed-form solution, the update for the cluster centers $\mathbf{c_{j}}$ becomes a Fréchet mean problem, which must be solved with an iterative Riemannian optimization routine:
$\mathbf{c_{j}}=\arg\min_{\mathbf{c}\in\mathcal{M}}\sum_{i=1}^{N}u_{ij}^{m}\delta_{\mathcal{M}}(\mathbf{x_{i}},\mathbf{c})^{2}.$ (105)
This approach is computationally expensive because it requires running a full Riemannian optimization for each cluster center in every iteration of the main algorithm. This leads to a time complexity of $\mathcal{O}(MN)$, where $M$ is the number of main clustering iterations and $N$ is the cost of a single Riemannian optimization.
RFK reformulates the objective function to avoid repeated optimization by expressing the optimal membership $u_{ij}$ as a direct function of the cluster centers $\{\mathbf{c_{j}}\}$:
$u_{ij}(\mathbf{C})=\left(\sum_{k=1}^{C}\left(\frac{\delta_{\mathcal{M}}(\mathbf{x_{i}},\mathbf{c_{j}})}{\delta_{\mathcal{M}}(\mathbf{x_{i}},\mathbf{c_{k}})}\right)^{\frac{2}{m-1}}\right)^{-1}.$ (106)
By substituting this expression back into the objective function, the problem is transformed into a single optimization problem that depends only on the cluster centers $\mathbf{C}$:
$J_{m}(\mathbf{C})$ $=\sum_{i=1}^{N}\sum_{j=1}^{C}\left(\left(\sum_{k=1}^{C}\left(\frac{\delta_{\mathcal{M}}(\mathbf{x_{i}},\mathbf{c_{j}})}{\delta_{\mathcal{M}}(\mathbf{x_{i}},\mathbf{c_{k}})}\right)^{\frac{2}{m-1}}\right)^{-1}\right)^{m}\delta_{\mathcal{M}}(\mathbf{x_{i}},\mathbf{c_{j}})^{2}$ (107)
$=\sum_{i=1}^{N}\left(\sum_{j=1}^{C}\delta_{\mathcal{M}}(\mathbf{x_{i}},\mathbf{c_{j}})^{\frac{2}{1-m}}\right)^{1-m}.$ (108)
The resulting objective function in Equation 108 is differentiable with respect to the cluster centers $\mathbf{c_{j}}$. It can therefore be minimized using a single, continuous Riemannian gradient-based optimization run, reducing the time complexity from $\mathcal{O}(MN)$ to $\mathcal{O}(N)$.
G.2 The Radan Optimizer
To efficiently solve the Riemannian optimization problem posed by the RFK algorithm, Yuan et al. (2025) introduce the Radan optimizer. Radan is a Riemannian adaptation of the Adaptive Nesterov Momentum (Adan) algorithm (Xie et al., 2024), an effective optimizer in Euclidean space.
Radan extends Adan to operate on Riemannian manifolds in the standard Riemannian optimization fashion: by integrating the exponential map and parallel transport. The update rules for Radan at iteration $t$ are as follows:
$\mathbf{m}_{t}$ $=\beta_{1,t}\varphi(\mathbf{m}_{t-1})+(1-\beta_{1,t})\mathbf{g}_{t}$ (109)
$\mathbf{v}_{t}$ $=\beta_{2,t}\varphi(\mathbf{v}_{t-1})+(1-\beta_{2,t})(\mathbf{g}_{t}-\varphi(\mathbf{g}_{t-1}))$ (110)
$\mathbf{z}_{t}$ $=\mathbf{g}_{t}+\beta_{2,t}(\mathbf{g}_{t}-\varphi(\mathbf{g}_{t-1}))$ (111)
$n_{t}$ $=\beta_{3}n_{t-1}+(1-\beta_{3})\|\mathbf{z}_{t}\|^{2}$ (112)
$\mathbf{u}_{t}$ $=\mathbf{m}_{t}+\beta_{2,t}\mathbf{v}_{t}$ (113)
$\alpha_{t}$ $=\frac{\eta_{t}}{\sqrt{n_{t}}+\epsilon}$ (114)
$\mathbf{x}_{t+1}$ $=\exp_{\mathbf{x}_{t}}(-\alpha_{t}\mathbf{u}_{t}),$ (115)
where:
  • $\mathbf{x}_{t}$ is the current point (parameter) on the manifold.
  • $\mathbf{g}_{t}$ is the Riemannian gradient of the loss function at $\mathbf{x}_{t}$.
  • $\varphi(\cdot)$ denotes parallel transport from tangent space $T_{\mathbf{x}_{t-1}}\mathcal{M}$ to $T_{\mathbf{x}_{t}}\mathcal{M}$.
  • $\mathbf{m}_{t}$ is the first-order moment (momentum) of the gradients.
  • $\mathbf{v}_{t}$ is the first-order moment of the gradient difference, accelerating convergence.
  • $n_{t}$ is the second-order moment for adaptive step sizing.
  • $\exp_{\mathbf{x}_{t}}(\cdot)$ is the exponential map, which ensures the new $\mathbf{x}_{t+1}$ remains on $\mathcal{M}$.
Appendix H Experiments
@@
H.1 Validating $\delta$-Hyperbolicity Estimation (Khrulkov et al. (2020))
@@
H.2 Validating Coordinate Learning (Gu et al 2018)